# Scaling of poroelastic coarsening and elastic arrest in crosslinked gels

### v1

@@ II.2 Generalizing LSW to coarsening of solvent-rich domains in gels:
The use of the diffusion length $\ell_{p}=\sqrt{D_{p}\ t}$ for the slow poroelastic limit is only appropriate if $\ell_{p}<\lambda(t)$; relaxation of the gel on scales larger than the domain size $\lambda(t)$ is not relevant for solvent injection into the gel by a shrinking domain of size $\lambda$. Thus, the slow poroelastic limit crosses over to the fast limit at a crossover domain size. Equating $\ell_{p}\sim\lambda(t)$ yields a time-independent crossover length
$\lambda_{\times}\sim\,\frac{\sigma}{G}\,\frac{\eta_{ps}}{\eta_{D}},$ (5)
@@
which magnifies the elastocapillary length $\sigma/G$ (which is of the order the mesh size of the gel) by a viscosity ratio. The crossover length can be large since the Darcy viscosity $\eta_{D}$ accounts for the solvent viscosity and friction with the instantaneous state of the polymers of the gel while the poroelastic viscosity in the viscoelastic regime of the gel, where coarsening takes place, accounts for polymer-polymer friction and local chain relaxation When the domain size $\lambda(t)<\lambda_{\times}$ the slow limit applies with coarsening scaling as $t^{1/4}$. In the opposite limit, the pressure gradients decay as $1/\lambda(t)$ which is appropriate in the fast poroelastic limit where the domain size scaling is $t^{1/3}$.
II.3 Elastic arrest of coarsening
@@
So far, we have focused on the capillary driving force for coarsening. At early times, the coarsening gel is in the viscoelastic regime, where polymers respond locally and transiently to the stress imposed by solvent expelled from shrinking domains. In this regime, elastic stresses relax on times shorter than the coarsening time and there is no spatially long-ranged and temporally persistent elastic response; those transient stresses therefore act primarily as a source of dissipation, slowing—but not halting—the capillarity-driven coarsening dynamics.
At sufficiently long times, $t>\tau_{\mathrm{el}}$, the local and transient viscoelastic polymer relaxation modes coupled to the coarsening-induced stress have relaxed, and the polymer network responds collectively and elastically [11, 22]. Although the elastic modulus is a property of the network at all times, it becomes dynamically operative as a long-ranged elastic stress that resists solvent flow only in the elastic regime, when stresses generated by solvent transport can no longer relax locally and a temporally permanent and a spatially long-ranged stress profile is established in the gel. In an elastic gel, capillary forces are balanced by gel stress at the domain boundary, determined self-consistently by elastic deformation of the entire surrounding medium. Elastic arrest is therefore most transparently understood as a force balance at the boundary of the shrinking domain: the Laplace capillary pressure ($\delta p=\sigma/\lambda)$ is canceled by the elastic stress of the gel.
@@
Crucially, the elastic stress at the domain boundary is not fixed a-priori. It must be determined by solving the elastic displacement problem in the gel surrounding the entire domain. For a spherical domain in an elastic solid [14], normal interfacial stress is of order
$\sigma_{\rm el}\;\sim\;G\,\frac{\delta\lambda}{\lambda}$ (6)
@@
where $G$ is the shear modulus of the gel and $\delta\lambda$ is the elastic displacement of the interface. Physically, $\delta\lambda$ represents the instantaneous radial deformation of the gel at the domain boundary induced by the capillary pressure acting on the curved interface; in the elastic regime it is defined at all times as the solution of the elastic equilibrium equations with the instantaneous interfacial pressure $\delta p(t)\sim\sigma/\lambda(t)$ applied as a boundary condition, which in general is time dependent because $\lambda(t)$ evolves during coarsening.
@@
Prior to arrest, the magnitude of $\delta\lambda$ is smaller than its long-time value because the long-ranged elastic stress field has not yet fully developed and therefore lags behind the capillary pressure. At long times, once the elastic stress has equilibrated spatially, the Laplace force balance at the domain boundary fixes $\delta\lambda$ by equating the elastic stress to the capillary pressure, $G(\delta\lambda/\lambda)\sim\gamma/\lambda$, which yields $\delta\lambda\sim\gamma/G$; thus, although the domain size $\lambda$ is mesoscopic, the interfacial displacement itself is set by a microscopic elastic length of order the gel mesh size. Because the elastic deformation extends over a region of size $\sim\lambda$, the domain size itself enters the solution: larger domains require elastic deformation over larger distances and therefore generate a different stress response. By contrast, the elastic stress at the solvent-rich domain boundary is not present in the viscoelastic, coarsening regime, where the unbalanced capillary pressure $\delta p$ is what drives solvent flow and coarsening.
@@
$\lambda_{\rm arrest}^{(\rm slow)}\sim\Bigg(\,\frac{\eta_{ps}\,\eta_{el}}{\eta_{D}^{2}}\,\frac{k_{0}\,\sigma^{2}}{G^{2}}\,\Bigg)^{1/4},$ (7)
@@
$\lambda_{\rm arrest}^{(\rm fast)}\sim\Bigg(\,\frac{\eta_{el}}{\eta_{D}}\,\frac{k_{0}\,\sigma}{G}\,\Bigg)^{1/3}.$ (8)
III Results
@@
In this section, we analyze the previous results to predict the scaling of the domain coarsening as a function of time and gel elastic modulus, as well as the modulus dependence of the arrest length. In addition, we show that this length can be magnified to a value far larger than the gel mesh size, by the viscosity ratios defined above. We present our analysis for two limiting cases of the dependence of the gel properties (permeability $k_{0}$ and interfacial tension with the solvent-rich phase $\sigma$) that enter the capillary-driven coarsening and eventual arrest. We keep the explicit dependence on the gel modulus $G$ since this is often changed and measured in experiments [9]. The first scenario we call the mesh-scaling gel where the polymer fraction in the gel is relatively small and the mesh size $\xi$ is the largest length in the homogeneous gel. The permeability in the Darcy equation has units of area and hence scales as $\xi^{2}$, while the tension (in units of $k_{B}T$) scales as $1/\xi^{2}$. The second scenario we call the melt-like gel where the polymer fraction in the gel is relatively large (order 1/2) and the mesh size is nanometric and can be of order the polymer persistence length and solvent molecule size. In contrast to the mesh-scaling gel with large pores which are the channels for solvent flow, the system may be conceptualized as somewhat swollen and crosslinked polymer melt. This is the limit that is applicable to the gels in [9] where the arrest length has been measured. For the melt-like gel, we keep $k_{0}$ and $\sigma$ as parameters which vary slowly with the gel mesh size, due to the involvement of the other length scales of the system.
@@
We first present the scaling relations for the coarsening dynamics and the arrest length as functions of the three viscosities defined above: The Darcy viscosity $\eta_{D}$, the poroelastic viscosity for stress diffusion in the viscoelastic regime of the polymers $\eta_{ps}$, and the viscosity $\eta_{el}=G\ \tau_{el}$ related to the time scale $\tau_{el}$, at which the viscoelastic regime crosses over to the elastic regime. Afterwards, we rewrite the scaling relations as functions of time scales associated with the Darcy and poroelastic viscosities, guided by the fact that a viscosity is the product of a modulus and a time scale. This effective modulus characterizes the short-time resistance of the polymer network to deformation by solvent flow. For scaling purposes, the distinction between this viscoelastic modulus and the long-time elastic modulus $G$ affects only prefactors and is therefore neglected since both are governed by the same crosslinked, network structure [23, 11]. We therefore write $\eta_{D}=G\ \tau_{D}$ and $\eta_{ps}=G\ \tau_{ps}$ where the dependence on $G$ reflects that the solvent flow, is impeded by friction with the gel even in its viscoelastic regime. The results for the scaling relations are shown in Tables 1 and 2.
@@
(i) Mesh-scaling gel: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim\left(\frac{k_{B}T}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
@@
(ii) Melt-like gel: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G}\,\frac{\eta_{ps}}{\eta_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
@@
(i) Mesh-scaling gel + $\eta\sim G\tau$: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\sigma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
@@
(ii) Melt-like gel + $\eta\sim G\tau$: $\sigma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\sigma}{G}\,\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\sigma}{G}\,\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\sigma^{2}}{G^{2}}\,\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
@@ IV Discussion
Melt-scaling gel: scaling of domain coarsening with time and elastic modulus: For this system, we keep the porosity and interfacial tension as parameters that vary only slowly with the modulus $G$, since for large polymer fraction gels (the melt-scaling limit), additional length scales related to the Kuhn length and solvent size can determine these parameters. However, despite this, the crosslinked, melt-scaling gel is still well characterized by its elastic modulus $G$ which remains the dominant parameter controlling both stress transmission and arrest. In this regime, fast poroelastic diffusion predicts coarsening which scales as $t^{1/3}$ while slow poroelastic diffusion coarsening scales as $t^{1/4}$. If we further take the viscosities to all scale as the product of the gel modulus $G$ and different effective relaxation times as in Table 2, for fast poroelastic coarsening, the dependence on the elastic modulus in all cases, for both the coarsening and the arrest lengths scale with $1/G^{1/3}$.
On the other hand, for slow poroelastic coarsening of the melt-like gel, we stress that both the coarsening and arrest lengths of Table 2 scale as $1/\sqrt{G}$. The arrest length is magnified from its microscopic or mesh-size value by the ratio of the time scales $\tau_{el}\gg\tau_{D}$ for fast coarsening and $\tau_{el}\tau_{ps}\gg\tau_{D}^{2}$ for slow coarsening. These large ratios reflect the separation between solvent transport and polymer stress relaxation processes: $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains activated by the solvent flow while $\tau_{D}$ is related to solvent flow in the presence of polymer friction in an essentially static configuration. The longest time scale is $\tau_{el}$ which is the crossover from viscoelasticity (activated by solvent flow) to elastic response when further coarsening is kinetically arrested by elastic stress acting over a region of order $\lambda$. We emphasize that the arrest discussed here refers to the breakdown of capillarity-driven, pressure-mediated coarsening; while other, much slower mechanisms of structural evolution cannot be excluded in principle, they lie outside the scope of the present theory.
@@
We now address the very slow drainage of solvent at time scales of about a day. For that process, also driven by capillarity involving the solvent in the gel and the excess solvent reservoir, the solvent flow is slow and we expect that the macroscopic drainage to depend on poroelastic diffusion, but with Darcy flow related to the solvent viscosity $\eta_{s}$ and not $\eta_{ps}$ since the polymers have relaxed all stresses on the macroscopic time scale. We thus estimate that the macroscopic time for drainage scales as $R^{2}/D_{D}$ where $R$ is the macroscopic sample size and the Darcy poroelastic diffusion constant $D_{D}$ scales as $k_{0}G/\eta_{s}$. For an order of magnitude numerical estimate, we take the permeability $k_{0}\sim\xi^{2}\sim(k_{B}\ T/G)^{2/3}$ and $\eta_{S}$ to be of the order of magnitude of the water viscosity, $10^{-3}\text{Pa~s}$. For $G\sim 100\text{kPa}$ as a typical rigidity [9], we find for a 1cm sample, a drainage time of the order of one day.

### v2

@@ II.2 Generalizing LSW to coarsening of solvent-rich domains in viscoelastic gels:
The use of the diffusion length $\ell_{p}=\sqrt{D_{p}\ t}$ for the slow poroelastic limit is only appropriate if $\ell_{p}<\lambda(t)$; relaxation of the gel on scales larger than the domain size $\lambda(t)$ is not relevant for solvent release n into the gel by a shrinking domain of size $\lambda$ (see Appendix C). Thus, the slow poroelastic limit crosses over to the fast limit at a crossover domain size. Equating $\ell_{p}\sim\lambda(t)$ yields a time-independent crossover length
$\lambda_{\times}\sim\,\frac{\gamma}{G}\,\frac{\eta_{ps}}{\eta_{D}},$ (5)
@@
which magnifies the elastocapillary length $\gamma/G$ (which is of the order the mesh size of the gel) by a viscosity ratio. The crossover length can be large since the Darcy viscosity $\eta_{D}$ accounts for the solvent viscosity and friction with the instantaneous state of the polymers of the gel while the poroelastic viscosity in the viscoelastic regime of the gel, where coarsening takes place, accounts for polymer-polymer friction and local chain relaxation When the domain size $\lambda(t)<\lambda_{\times}$ the slow limit applies with coarsening scaling as $t^{1/4}$. In the opposite limit, the pressure gradients decay as $1/\lambda(t)$ which is appropriate in the fast poroelastic limit where the domain size scaling is $t^{1/3}$.
III Elastic arrest
@@
So far, we have focused on the capillary driving force for coarsening. At early times, the coarsening gel is in the viscoelastic regime, where polymers respond locally and transiently to the stresses generated by capillary tractions (normal stresses) at curved solvent–gel interfaces. In this regime, there is no spatially long-ranged and temporally persistent elastic response; transient polymer stresses therefore act primarily as a source of dissipation, slowing—but not halting—the capillarity-driven coarsening dynamics. Mechanically, the Young–Laplace condition prescribes a normal traction (normal stress) at the domain boundary; because a long-ranged, persistent deviatoric stress field is not possible in the viscoelastic regime, this traction is transmitted primarily through an isotropic pressure perturbation in the solvent pores (i.e., the pore-pressure) of the gel. Matching this perturbation to the far-field pressure necessarily generates a nonzero pressure gradient, which by Darcy’s law drives solvent flow and hence coarsening.
However, at sufficiently long times, $t>\tau_{\mathrm{el}}$, the local and transient viscoelastic polymer relaxation modes coupled to the coarsening-induced stress have relaxed, and the polymer network responds collectively and elastically [11, 22]. Although the elastic modulus is a property of the network at all times, the gel stress is only spatially long-ranged and temporally persistent in this elastic regime. In this regime, the Young–Laplace interfacial traction (i.e., the normal force per unit area—equivalently, the normal component of the stress—transmitted across the interface) has a characteristic magnitude set by the Laplace-pressure scale, $\delta p\sim\gamma/\lambda$, at the solvent–gel interface. As shown in Appendix D, boundary traction is then balanced by the long-ranged deviatoric (non-isotropic) stress field established in the gel, but the pressure gradient is zero.
Crucially, the resulting pore-pressure field is not assumed to have zero gradient: it is determined by solving the elastic equilibrium problem in the surrounding gel. Writing the gel stress as $\bm{\sigma}=-p\,\mathbf{I}+2G\bm{\varepsilon}$ and imposing spherical mechanical equilibrium, the elastic displacement/strain field that balances the interfacial traction yields $dp/dr=0$, i.e. the isotropic pore pressure outside the domain is spatially uniform (with the constant fixed by the far-field boundary condition) (see Appendix D). Therefore, in this elastic regime the pressure gradient entering Darcy’s law vanishes, so the pressure-gradient mechanism for solvent transport is eliminated. By Darcy’s law, this removes the driving force for solvent flow and thus halts the capillarity-driven coarsening. We therefore refer to elastic arrest as the regime in which the interfacial capillary traction is balanced by the gel’s elastic shear stress, leaving no pore-pressure gradient to drive further Darcy permeation.
@@
It is also important to note that the elastic stress at the domain boundary is not fixed a-priori. It must be determined by solving the elastic displacement problem in the gel surrounding the entire domain. For a spherical domain embedded in an infinite elastic gel [14], the radial displacement field decays as (see Appendix D)
$u(r)\sim\delta\lambda\left(\frac{\lambda}{r}\right)^{2},$ (6)
so that the normal stress at the interface is proportional to the interfacial strain,
$\sigma_{\rm el}\sim G\frac{\delta\lambda}{\lambda},$ (7)
@@
up to a numerical prefactor. Physically, $\delta\lambda$ represents the instantaneous radial deformation of the gel at the domain boundary induced by the capillary pressure acting on the curved interface. It is not an imposed equilibrium mismatch between two stress-bearing phases; rather, it arises dynamically from the requirement that the gel satisfy mechanical equilibrium with the instantaneous Laplace pressure $\delta p(t)\sim\gamma/\lambda(t)$. In the elastic regime it is defined at all times as the solution of the elastic equilibrium equations with the instantaneous interfacial pressure applied as a boundary condition, which in general is time dependent because $\lambda(t)$ evolves during coarsening.
@@
At long times, once the elastic stress has equilibrated spatially, the Laplace force balance at the domain boundary fixes $\delta\lambda$ by equating the elastic stress to the capillary pressure, $G(\delta\lambda/\lambda)\sim\gamma/\lambda$, which yields $\delta\lambda\sim\gamma/G$; thus, although the domain size $\lambda$ is mesoscopic, the interfacial displacement itself is set by a microscopic elastic length of order the gel mesh size. Importantly, this boundary force balance fixes the elastic displacement $\delta\lambda$ but it does not by itself determine the arrested domain size itself $\lambda_{\rm arrest}$ which is determined kinetically as discussed below.
To summarize: In the elastic regime of the gel there is only a deviatoric (non-isotropic) stress induced by the capillary traction at the boundary of the solvent-rich domain and the pressure gradient is zero (see Appendix D). Since $\dot{\lambda}\sim-(k_{0}/\eta_{D})\nabla p$, the solvent flux ceases and coarsening stops. By contrast, in the viscoelastic coarsening regime, the long-ranged elastic stress field has not yet developed, so the capillary pressure remains effectively unbalanced, a pressure gradient exists in the gel and by the Darcy flow equation solvent flow and coarsening proceed. Such elastic arrest does not occur in a coherent AABB solid solution (of two elastic phases), since mass transport is diffusive and driven by gradients of chemical potential, which is not proportional to the pressure gradient in an elastic system. The chemical potential contain contributions from the entire elastic stress field (both isotropic pressure and deviatoric, non-isotropic stress). Consequently, in the gel elastic arrest of the solvent-rich, fluid domains occurs when $\nabla p=0$, whereas in a solid solution elasticity modifies the thermodynamic driving force itself rather than eliminating transport through the vanishing of a pressure gradient.
@@
$\lambda_{\rm arrest}^{(\rm slow)}\sim\Bigg(\,\frac{\eta_{ps}\,\eta_{el}}{\eta_{D}^{2}}\,\frac{k_{0}\,\gamma^{2}}{G^{2}}\,\Bigg)^{1/4},$ (8)
@@
$\lambda_{\rm arrest}^{(\rm fast)}\sim\Bigg(\,\frac{\eta_{el}}{\eta_{D}}\,\frac{k_{0}\,\gamma}{G}\,\Bigg)^{1/3}.$ (9)
IV Results
@@
In this section, we analyze the previous results to predict the scaling of the domain coarsening as a function of time and gel elastic modulus, as well as the modulus dependence of the arrest length. In addition, we show that this length can be magnified to a value far larger than the gel mesh size, by the viscosity ratios defined above. We present our analysis for two limiting cases of the dependence of the gel properties (permeability $k_{0}$ and interfacial tension with the solvent-rich phase $\gamma$) that enter the capillary-driven coarsening and eventual arrest. We keep the explicit dependence on the gel modulus $G$ since this is often varied and measured in experiments [9]. The first scenario we call the mesh-scaling gel where the polymer fraction in the gel is relatively small and the mesh size $\xi$ is the largest length in the homogeneous gel. The permeability in the Darcy equation has units of area and hence scales as $\xi^{2}$, while the tension (in units of $k_{B}T$) scales as $1/\xi^{2}$. The second scenario we call the melt-like gel where the polymer fraction in the gel is relatively large (order 1/2) and the mesh size is nanometric and can be of order the polymer persistence length and solvent molecule size. In contrast to the mesh-scaling gel with large pores which are the channels for solvent flow, the system may be conceptualized as somewhat swollen and crosslinked polymer melt. This is the limit that is applicable to the gels in [9] where the arrest length has been measured. For the melt-like gel, we keep $k_{0}$ and $\gamma$ as parameters which vary slowly with the gel mesh size, due to the involvement of the other length scales of the system.
@@
We first present the scaling relations for the coarsening dynamics and the arrest length as functions of the three viscosities defined above: The Darcy viscosity $\eta_{D}$, the poroelastic viscosity for stress diffusion in the viscoelastic regime of the polymers $\eta_{ps}$, and the viscosity $\eta_{el}=G\ \tau_{el}$ related to the time scale $\tau_{el}$, at which the viscoelastic regime (present at short times even in a crosslinked gel) crosses over to the elastic regime. Afterwards, we rewrite the scaling relations as functions of time scales associated with the Darcy and poroelastic viscosities, guided by the fact that a viscosity is the product of a modulus and a time scale. This effective modulus characterizes the short-time resistance of the polymer network to deformation by solvent flow. For scaling purposes, the distinction between this viscoelastic modulus and the long-time elastic modulus $G$ affects only prefactors and is therefore neglected since both are governed by the same crosslinked, network structure [22, 11]. We therefore write $\eta_{D}=G\ \tau_{D}$ and $\eta_{ps}=G\ \tau_{ps}$ where the dependence on $G$ reflects that the solvent flow, is impeded by friction with the gel even in its viscoelastic regime. The results for the scaling relations are shown in Tables 1 and 2.
@@
(i) Mesh-scaling gel: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\gamma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim\left(\frac{k_{B}T}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
@@
(ii) Melt-like gel: $\gamma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\gamma}{\eta_{D}}\,t\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\gamma}{G}\,\frac{\eta_{\rm el}}{\eta_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\gamma^{2}}{G}\,\frac{\eta_{ps}}{\eta_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\eta_{ps}\eta_{\rm el}}{\eta_{D}^{2}}\right)^{1/4}$
@@
(i) Mesh-scaling gel + $\eta\sim G\tau$: $G\sim k_{B}T/\xi^{3}$, $k_{0}\sim\xi^{2}$, $\gamma\sim k_{B}T/\xi^{2}$ Fast $\lambda(t)\sim(k_{B}T)^{1/3}G^{-1/3}\left(\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{B}T}{G}\right)^{1/3}\left(\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
@@
(ii) Melt-like gel + $\eta\sim G\tau$: $\gamma,k_{0}$ independent Fast $\lambda(t)\sim\left(\frac{k_{0}\gamma}{G}\,\frac{t}{\tau_{D}}\right)^{1/3}$ $\lambda_{\rm arrest}^{(\mathrm{fast})}\sim\left(\frac{k_{0}\gamma}{G}\,\frac{\tau_{el}}{\tau_{D}}\right)^{1/3}$
Slow $\lambda(t)\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\tau_{ps}}{\tau_{D}^{2}}\,t\right)^{1/4}$ $\lambda_{\rm arrest}^{(\mathrm{slow})}\sim\left(\frac{k_{0}\gamma^{2}}{G^{2}}\,\frac{\tau_{ps}\tau_{el}}{\tau_{D}^{2}}\right)^{1/4}$
@@ V Discussion
Melt-like gel: scaling of domain coarsening with time and elastic modulus: For this system, we keep the porosity and interfacial tension as parameters that vary only slowly with the modulus $G$, since for large polymer fraction gels (the melt-like limit), additional length scales related to the Kuhn length and solvent size can determine these parameters. However, despite this, the crosslinked, melt-like gel is still well characterized by its elastic modulus $G$ which remains the dominant parameter controlling both stress transmission and arrest. In this regime, fast poroelastic diffusion predicts coarsening which scales as $t^{1/3}$ while slow poroelastic diffusion coarsening scales as $t^{1/4}$. If we further take the viscosities to all scale as the product of the gel modulus $G$ and different effective relaxation times as in Table 2, for fast poroelastic coarsening, the dependence on the elastic modulus in all cases, for both the coarsening and the arrest lengths scale with $1/G^{1/3}$.
On the other hand, for slow poroelastic coarsening of the melt-like gel, we stress that both the coarsening and arrest lengths of Table 2 scale as $1/\sqrt{G}$. The arrest length is magnified from its microscopic or mesh-size value by the ratio of the time scales $\tau_{el}\gg\tau_{D}$ for fast coarsening and $\tau_{el}\tau_{ps}\gg\tau_{D}^{2}$ for slow coarsening. These large ratios reflect the separation between solvent transport and polymer stress relaxation processes: $\tau_{ps}$ is related to the local and transient viscoelastic relaxation of the polymer chains activated by the solvent flow, while $\tau_{D}$ is related to solvent flow in the presence of polymer friction in an essentially static configuration. The longest time scale is $\tau_{el}$ which is the crossover from viscoelasticity (activated by solvent flow) to elastic response when further coarsening is kinetically arrested by elastic stress acting over a region of order $\lambda$. We emphasize that the arrest discussed here refers to the breakdown of capillarity-driven, pressure-mediated coarsening; while other, much slower mechanisms of structural evolution cannot be excluded in principle, they lie outside the scope of the present theory.
@@
We now address the very slow drainage of solvent at time scales of many hours. That process, is also driven by capillarity involving the solvent in the gel near its boundary with an excess solvent reservoir. The solvent flow is slow and we expect that the drainage initiated at the sample boundaries (not included in the other sections of this paper) will depend on poroelastic diffusion in that region, with the Darcy flow near the boundary related to the solvent viscosity $\eta_{s}$ and not $\eta_{ps}$; this is because the polymers have relaxed all stresses on the macroscopic time scale. We thus estimate that the macroscopic time for drainage at the boundary scales as $R^{2}/D_{D}$ where $R$ is a macroscopic size and the Darcy poroelastic diffusion constant $D_{D}$ scales as $k_{0}G/\eta_{s}$. For an order of magnitude numerical estimate, we take the permeability $k_{0}\sim\xi^{2}\sim(k_{B}\ T/G)^{2/3}$ and $\eta_{S}$ to be of the order of magnitude of the water viscosity, $10^{-3}\text{Pa~s}$. For $G\sim 100\text{kPa}$ as a typical rigidity [9], we find for a 1cm sample, a drainage time of the order of one day.